# Label Dropout: Improved Deep Learning Echocardiography Segmentation Using Multiple Datasets With Domain Shift and Partial Labelling

### v1

@@ 1 Introduction
Training naively with partially labelled datasets such as these can cause a conflict in the supervision due to structures being labelled as foreground in some samples and background in others. A number of methods have been proposed to deal with such data by modifying the loss function to deal with the conflict. For example, Petit et al [9] proposed a method that altered the loss function from categorical cross entropy (CCE) to a summation of binary cross entropy (BCE) losses, one for each foreground class. For samples with missing ground truth labels, only the BCE terms with ground truth labels were computed. A related approach was proposed by Shi et al [12]. Their work proposed a marginal loss, in which any missing label was merged with the background class. In these merged regions, the loss function considered both the missing label and the background to be a correct prediction. Finally, Mariscal-Harana et al [6] proposed the adaptive loss, which was similar in concept to the approach of Petit et al [9], but removed the loss for the label not present in the ground truth from the CCE loss calculation.
@@ 2.0.3 Label dropout:
As we will show in Section 3, a significant drop in performance of the adaptive loss model was observed when training with partially labelled datasets which have a domain shift between them. It was hypothesised that this was due to the model learning to associate domain specific characteristics with the presence of labels. Therefore, we propose the label dropout scheme, which aims to break the link between domain characteristics and the presence or absence of certain labels. In label dropout, we introduce a random probability of a label being removed (i.e. set to background) from the ground truth mask of each sample during training.
@@ 3 Experiments and Results
3.0.1 Experiment 1 - The need for multiple diverse training datasets:
@@ 3.0.4 Experiment 4 - Label dropout:
In Fig. 4, sample model predictions from this experiment are also displayed. When there is no label dropout, the LA in the Unity Imaging dataset has a visibly worse segmentation compared to the adaptive loss model with 50% label dropout. This experiment shows that label dropout can improve model performance when training with diverse partially labelled datasets. The Dice scores for the LV and LVM were similar for both models for each dataset.
@@
To quantify these improvements, the LVM and LA were manually segmented by a trained observer (and checked by a cardiologist) in a random sample of 20 images from EchoNet Dynamic test set. The mean LVM Dice scores achieved for the adaptive loss with and without label dropout were 0.319 and 0.518 respectively, and for LA they were 0.553 and 0.689 respectively. These respresent Dice increases of 0.199 and 0.136 (or 62% and 25%). Model predictions for 10 random images from these 20 are displayed in Supplementary Fig. 1 with a summary of the Dice scores in Supplementary Table 1.

### v2

@@ 1 Introduction
Training naively with partially labelled datasets such as these can cause a conflict in the supervision due to structures being labelled as foreground in some samples and background in others. A number of methods have been proposed to deal with such data by modifying the loss function to deal with the conflict. For example, Petit et al [petit2018handling] proposed a loss function that summed the binary cross entropy (BCE) losses for each foreground class. For samples with missing ground truth labels, only the BCE terms with ground truth labels were computed. A related approach was proposed by Shi et al [12]. Their work proposed a marginal loss, which was a modification of the categorical cross entropy (CCE) loss function, in which any missing label was merged with the background class. In these merged regions, the loss function essentially considered both the missing label and the background to be a correct prediction. Finally, Mariscal-Harana et al [6] proposed the adaptive loss, which was similar in concept to the approach of Petit et al [petit2018handling], but removed the loss for the label not present in the ground truth from the CCE loss calculation.
@@ 2.0.3 Label dropout:
As we will show in Section 3, a significant drop in performance of the adaptive loss model was observed when training with partially labelled datasets which have a domain shift between them. It was hypothesised that this was due to the model learning to associate domain specific characteristics with the presence of labels. Therefore, we propose the label dropout scheme, which aims to break the link between domain characteristics and the presence or absence of certain labels. In label dropout, we introduce a probability of a label being removed (i.e. set to background) from the ground truth mask of each sample during training. Each training sample is considered individually on the fly during training. Label dropout is applied only to partially labelled classes (i.e. those that are missing in some ground truth masks). In the case of there being multiple missing partially labelled classes, only one of them is dropped out, with an equal chance of each of them being chosen.
@@ 3 Experiments and Results
3.0.1 Experiment 1 - The need for training with multiple diverse datasets:
@@ 3.0.4 Experiment 4 - Label dropout:
In Fig. 4, sample model predictions from this experiment are also displayed. When there is no label dropout, the LA in the Unity Imaging dataset has a visibly worse segmentation compared to the adaptive loss model with 50% label dropout. This experiment shows that label dropout can improve model performance when training with diverse partially labelled datasets and a Wilcoxon Signed Rank Test confirms this with a p-value of <0.01. The Dice scores for the LV and LVM were similar for both models for each dataset.
@@
To quantify these improvements, the LVM and LA were manually segmented by a trained observer (and checked by a cardiologist) in a random sample of 20 images from EchoNet Dynamic test set and some key results are shown in Table 2. The LVM and LA Dice scores represent Dice increases of 0.199 and 0.136 (or 62% and 25%) respectively, supported by a Wilcoxon Signed Rank Test with a p-value of <0.05. Model predictions for 10 random images from these 20 are displayed in Supplementary Fig. 1.
Table 2: Mean test Dice scores over 20 random images from EchoNet Dynamic when trained on EchoNet Dynamic, CAMUS and Unity Imaging data. Dice scores are displayed for each of the three cardiac structures segmented by the adaptive loss model with and without label dropout. Aug = augmentation, LD = label dropout, LV = left ventricle, LVM = left ventricular myocardium, LA = left atrium.
LV LVM LA
Adaptive (w/ aug) 0.892 0.319 0.553
Adaptive (w/ aug & LD) 0.910 0.518 0.689